OpenStudy (hungergames13):

Which ordered pair is the solution to the system of equations? x-4y=6 3x+y=-8 A. (2, −1) B. (1, −9) C. (−2, −2) D. (−6, −3)

OpenStudy (gabebae):

Hello

OpenStudy (hungergames13):

hi

OpenStudy (gabebae):

substitute the x and y values for all 4 choices and see which one is correct

OpenStudy (ihelpyoulearn):

@math&ing001

OpenStudy (math&ing001):

Or you can solve the system: -3*(x-4y=6) ie. -3x+12y=-18 +3x+ y =-8 13y= -26 y=-2 Now find x by replacing y with its value in one of the equations, then you'll get your ordered pair (x,y)

OpenStudy (hungergames13):

oh ok

OpenStudy (math&ing001):

So what did you get for x ?

OpenStudy (hungergames13):

hold on

OpenStudy (hungergames13):

i think its D

OpenStudy (math&ing001):

Not quite, did you try to get x the way I told you ?

OpenStudy (hungergames13):

umm hold on let me try it again

OpenStudy (hungergames13):

is it A?

OpenStudy (math&ing001):

I'm not interested in the answer as much as I'm interested in the way you find it. Please explain how you did to find those answers.

OpenStudy (hungergames13):

well I found A by doing\[2-4*-1\]

OpenStudy (hungergames13):

and it gave me 6

OpenStudy (math&ing001):

Yes but did you try for the second equation ? It has to be correct for both !

OpenStudy (hungergames13):

ohhh its c

OpenStudy (math&ing001):

Correct !

OpenStudy (hungergames13):

yay:D